How many gallons of an 80% alcohol solution must be mixed with 30 gallons of a 22% solution to obtain a solution that is 70% alcohol?
Answer by checkley71(8403)

.8X+.22*30=.7(X+30)
.8X+6.6=.7X+21
.8X-.7X=21-6.6
.1X=14.4
X=14.4/.1
X=144 GALLONS OF 80% SOLUTION IS REQUIRED.
PROOF
.8*144+.22*30=.7(144+30)
115.2+6.6=121.8
121.8=121.8